# Well-posedness of a first-order-in-time model for nonlinear acoustics with nonhomogeneous boundary conditions

### v2

@@ 2.1.1 Galerkin Approximation
To show existence of a solution, we use Galerkin’s method. That is, we use eigenvectors of $\mathcal{A}$ to construct a sequence $\{u_{m}\}_{m\in\mathbb{N}}$ by using linear combinations of $m$ eigenvectors. The construction of this sequence is well posed in the usual sense with absolutely continuous coefficients in time. It is determined by solving an ODE, obtained by projecting (2) down to the span of the first $m$ eigenvectors of $\mathcal{A}$. We omit the details here and refer to e.g. [5, Chapter 7] and [11, Section 2.3.1].
@@ 2.1.2 Energy estimates
Testing equation (2) with $\phi(s):=c_{\mathcal{A}}^{2}C_{\mathscr{A}}^{2}u_{m}(s)+\mathcal{A}^{\sigma}u_{m}(s)$, where $\sigma,c_{\mathcal{A}},C_{\mathscr{A}}$ as in Assumption 2.1, gives
$\langle\partial_{t}u_{m}(s)+\mathfrak{A}u_{m}(s)+\mathfrak{B}[u_{m}(s),{u_{*}}(s)]+\mathfrak{B}[{u_{*}}(s),u_{m}(s)],\phi(s)\rangle=\langle f(s),\phi(s)\rangle.$
Using assumption 2.1(ii) that $\mathcal{A}$ and its fractional powers are self adjoint and $\mathscr{A}$ is skew symmetric (2.1(ii)), while keeping in mind that $u_{m}(s)\in{D}(\mathcal{A})$ is absolutely continuous with respect to time, equation (2.1.2) implies
@@
where we omit the time argument $s$ of $u_{m}(s)$ in the following for reasons of readability. Using boundedness of $\mathscr{A}$, Young’s inequality and (2.1) with $\sigma\leq\frac{\sigma+1}{2}$ we get
$|\langle\mathscr{A}u_{m},\mathcal{A}^{\sigma}u_{m}\rangle|\leq\|\mathscr{A}u_{m}\|_{H}\|\mathcal{A}^{\sigma}u_{m}\|_{H}\leq\frac{c_{\mathcal{A}}^{2}C^{2}_{\mathscr{A}}}{2}\|\mathcal{A}^{\frac{1}{2}}u_{m}\|_{H}^{2}+\frac{1}{2}\|\mathcal{A}^{\frac{1+\sigma}{2}}u_{m}\|^{2}_{H},$
[…]
@@
Using (2.1) with $\frac{1-\sigma}{2}\leq\frac{1+\sigma}{2}$ and Young’s inequality twice we get
@@
Using the estimate (2.1(iii)) on $\mathfrak{B}$, Young’s inequality, the embedding (2.1) $X\hookrightarrow C([0,T];W)$ and (2.1) with $\frac{1-\sigma}{2}\leq\frac{\sigma+1}{2}$ we get
$\begin{split}2&\int_{0}^{t}|\langle\mathfrak{B}[u_{m},{u_{*}}]+\mathfrak{B}[{u_{*}},u_{m}],\phi\rangle|ds\\ &\leq 2\int_{0}^{t}C_{\mathfrak{B}}(\|\mathcal{A}^{\frac{1+\sigma}{2}}u_{m}\|_{H}\|\mathcal{A}^{\frac{\sigma}{2}}{u_{*}}\|_{H}+\|\mathcal{A}^{\frac{1+\sigma}{2}}{u_{*}}\|_{H}\|\mathcal{A}^{\frac{\sigma}{2}}u_{m}\|_{H})\|\mathcal{A}^{\frac{1-\sigma}{2}}\phi\|_{H}\\ &\leq\int_{0}^{t}\frac{C_{\mathfrak{B}}^{2}}{\|{u_{*}}\|_{X}}(\|\mathcal{A}^{\frac{1+\sigma}{2}}u_{m}\|_{H}^{2}\|\mathcal{A}^{\frac{\sigma}{2}}{u_{*}}\|_{H}^{2}+\|\mathcal{A}^{\frac{1+\sigma}{2}}{u_{*}}\|_{H}^{2}\|\mathcal{A}^{\frac{\sigma}{2}}u_{m}\|_{H}^{2})+{\|{u_{*}}\|_{X}}\|\mathcal{A}^{\frac{1-\sigma}{2}}\phi\|_{H}^{2}ds\\ &\leq{C_{\mathfrak{B}}^{2}}\|{u_{*}}\|_{X}(C_{T}^{2}\int_{0}^{t}\|\mathcal{A}^{\frac{1+\sigma}{2}}u_{m}\|_{H}^{2}ds+\max_{s\in[0,t]}\|\mathcal{A}^{\frac{\sigma}{2}}u_{m}\|_{H}^{2})\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad+(c_{\mathcal{A}}C_{A}+1)^{2}{\|{u_{*}}\|_{X}}\int_{0}^{t}\|\mathcal{A}^{\frac{1+\sigma}{2}}u_{m}\|_{H}^{2}ds\\ &\leq\|{u_{*}}\|_{X}(\max\{C_{T}^{2},1\}C_{\mathfrak{B}}^{2}+(c_{\mathcal{A}}C_{A}+1)^{2})\mathcal{E}[\mathcal{A}^{\frac{\sigma}{2}}u_{m}](t)\end{split}$
@@
${C_{A}}\mathcal{E}[u_{m}](t)+\frac{1}{4}\mathcal{E}[\mathcal{A}^{\frac{\sigma}{2}}u_{m}](t)\leq\delta^{-1}\|f\|_{Y}^{2}+(1+C_{A}c_{\mathcal{A}}^{2})\|u_{m}(0)\|^{2}_{W}.$
In particular, due to (2.1.2), $\|u_{m}\|^{2}_{L^{2}(0,T;U)}\leq C_{U}(\|f\|_{Y}^{2}+\|u_{m}(0)\|^{2}_{W})$ with $C_{U}=4\max\{\delta^{-1},(1+C_{A}c_{\mathcal{A}}^{2})\}$ and to estimate $\partial_{t}u_{m}$ we take the $V^{*}$ norm in equation (2) using (2.1(iii)), (2.1) to obtain
$\begin{split}\|\partial_{t}u_{m}\|_{V^{*}}&\leq\|\mathfrak{A}u_{m}\|_{V^{*}}+\|\mathfrak{B}[u,{u_{*}}]\|_{V^{*}}+\|\mathfrak{B}[{u_{*}},u]\|_{V^{*}}+\|f\|_{V^{*}}\\ &\leq C_{\mathfrak{A}}\|u_{m}\|_{U}+C_{\mathfrak{B}}\|{u_{*}}\|_{X}(\|u_{m}\|_{U}+\|u_{m}\|_{W})+\|f\|_{V^{*}}\\ &\leq C_{V}\|u_{m}\|_{U}+\|f\|_{V^{*}}.\end{split}$
@@
$\|\partial_{t}u_{m}\|_{L^{2}(0,T;V^{*})}^{2}\leq 2C_{U}C_{V}^{2}(\|f\|_{Y}^{2}+\|u_{m}(0)\|^{2}_{W})+\|f\|_{Y}^{2}=2(C_{U}C_{V}^{2}+1)(\|f\|_{Y}^{2}+\|u_{m}(0)\|^{2}_{W}).$
Thus, combing this with (2.1.2) we obtain the a priori estimate
$\|u_{m}\|_{X}\leq C_{G}(\|f\|_{Y}+\|u_{m}(0)\|_{W})$
@@ 2.1.3 Existence and uniqueness
It is therefore left to show that $u(0)=g$ holds, which we do by closely following the argument in [5, Section 7.1.2 c]. First of all, since $X\hookrightarrow C([0,T];W)$, we know that imposing an initial condition as $u(0)$ is sensible. Then by partial integration with respect to time in (2.1.3), we have on the one hand for some $v\in C^{1}([0,T];W)\subset L^{2}(0,T;V)$ with $v(T)=0$
$\langle\partial_{t}v,u\rangle_{Y}+\langle\mathfrak{A}u+\mathfrak{B}[u,{u_{*}}]+\mathfrak{B}[{u_{*}},u],v\rangle_{Y}=\langle f,v\rangle_{Y}+\langle u(0),v(0)\rangle_{V^{*}}$
@@
$\lim_{n\to\infty}-\langle\partial_{t}v,u_{m}\rangle_{Y}+\langle\mathfrak{A}u_{m}+\mathfrak{B}[u_{m},{u_{*}}]+\mathfrak{B}[{u_{*}},u_{m}],v\rangle_{Y}=\langle f,v\rangle_{Y}+\langle g,v(0)\rangle_{V^{*}}$
since $u_{m}(0)\to g$ weakly in $V$. Comparing both expression we see that due to arbitrariness of $v(0)$, we obtain $u(0)=g$. Uniqueness of $u$ follows by standard arguments, due to the linearity of (2), see [11, Section 3.2.3] for details. ∎
@@ 2.2 Contraction argument for the bilinear problem
$\|\mathcal{F}^{\prime}_{u_{1}}-\mathcal{F}^{\prime}_{u_{2}}\|_{\mathcal{L}(X,Y)}\leq K\|u_{1}-u_{2}\|_{X},\quad u_{1},\,u_{2}\in D_{*}.$
@@
A proof of Lemma 2.1 can be found in e.g. [12, 13].
@@
Let $T>0$ and Assumption 2.1 hold with $\sigma\in[0,1]$. Then, for small enough $r>0$ and any $(f,g)\in L^{2}(0,T;D(\mathcal{A}^{\frac{1-\sigma}{2}})^{*})\times{D}(\mathcal{A}^{\frac{\sigma}{2}})$ such that
@@
The proof follows a similar argumentation as the proof of [11, Theorem 3.2]. For convenience of the reader, we restate it here. As in Theorem 2.1, we use the short notations $Y:=L^{2}(0,T;D(\mathcal{A}^{\frac{1-\sigma}{2}})^{*})$, $U:=D(\mathcal{A}^{\frac{1+\sigma}{2}})$, $V:=D(\mathcal{A}^{\frac{1-\sigma}{2}})$ and $W:={D}(\mathcal{A}^{\frac{\sigma}{2}})$
@@
$\mathcal{F}u:=(\partial_{t}u+\mathfrak{A}u+\mathfrak{B}[u,u],u(0))-(f,g).$
@@
Then, by assumptions 2.1(iii) (2.1(iii)) and (2.1) that yield boundedness of $\mathfrak{B}:X\times X\to Y$ we have
$\frac{\|\mathfrak{B}[h,h]\|_{Y}}{\|h\|_{X}}\leq\frac{(C_{\mathfrak{B}}^{2}C_{T}^{2}\|h\|_{X}^{2}\int_{0}^{T}\|h\|_{U}^{2}ds)^{\frac{1}{2}}}{\|h\|_{X}}\leq{C_{\mathfrak{B}}}C_{T}\|h\|_{X}\to 0\quad(\|h\|_{X}\to 0)$
@@
$\mathcal{F}^{\prime}_{u_{*}}:\xi\mapsto(\partial_{t}\xi+\mathfrak{A}\xi,\xi(0))+(\mathfrak{B}[\xi,{u_{*}}]+\mathfrak{B}[u_{*},\xi],0),$
@@
$\|(\mathcal{F}^{\prime}_{u_{1}}-\mathcal{F}^{\prime}_{u_{2}})u\|_{Y}\leq\|\mathfrak{B}[u,u_{1}-u_{2}]+\mathfrak{B}[u_{1}-u_{2},u]\|_{Y}\leq 2C_{\mathfrak{B}}C_{T}\|u_{1}-u_{2}\|_{X}\|u\|_{X}$
@@
$\beta K\|-(\partial_{t}\cdot+\mathfrak{A}\cdot,\cdot(0))^{-1}(f,g)\|_{X}\leq\frac{1}{2}.$

### v3

@@ 2.1.1 Galerkin Approximation
To show existence of a solution, we use Galerkin’s method. That is, we approximate solutions to (2) by $u_{m}(t)=\sum_{k=1}^{m}\xi_{k}^{m}(t)\phi_{k}\in U_{m}=\text{span}\{\phi_{k}:k=1,\dots,m\}\subset U$ satisfying
$\langle\partial_{t}u_{m}+\mathfrak{A}u_{m}+\mathfrak{B}[u_{m},{u_{*}}]+\mathfrak{B}[{u_{*}},u_{m}],\phi\rangle=\langle f,\phi\rangle,\quad u_{m}(0)=g_{m}$ (19)
for all $\phi\in U_{m}\subset V$. Here $g_{m}$ is a projection of $g$ onto $U_{m}$. The construction of the sequence $\{u_{m}\}_{m\in\mathbb{N}}$ is well posed in the usual sense with absolutely continuous coefficients $\xi_{k}^{m}$. We omit the details here and refer to e.g. [Evans:2010, Chapter 7] and [QuadraticWave, Section 2.3.1].
@@ 2.1.2 Energy estimates
Testing equation (19) with $\phi(s):=c_{\mathcal{A}}^{2}C_{\mathscr{A}}^{2}u_{m}(s)+\mathcal{A}^{\sigma}u_{m}(s){\color[rgb]{0,0,0}\,\,\in U_{m}}$, where $\sigma,c_{\mathcal{A}},C_{\mathscr{A}}$ as in Assumption 2.1, gives
$\langle\partial_{t}u_{m}(s)+\mathfrak{A}u_{m}(s)+\mathfrak{B}[u_{m}(s),{u_{*}}(s)]+\mathfrak{B}[{u_{*}}(s),u_{m}(s)],\phi(s)\rangle=\langle f(s),\phi(s)\rangle.$ (20)
Using assumption 2.1(ii) that $\mathcal{A}$ and its fractional powers are self adjoint and $\mathscr{A}$ is skew symmetric (10), while keeping in mind that $u_{m}(s)\in{D}(\mathcal{A})$ is absolutely continuous with respect to time, equation (20) implies
@@
where we omit the time argument $s$ of $u_{m}(s)$ in the following for reasons of readability. Using boundedness of $\mathscr{A}$, Young’s inequality and (14) with $\sigma\leq\frac{\sigma+1}{2}$ we get
$|\langle\mathscr{A}u_{m},\mathcal{A}^{\sigma}u_{m}\rangle|\leq\|\mathscr{A}u_{m}\|_{H}\|\mathcal{A}^{\sigma}u_{m}\|_{H}\leq\frac{c_{\mathcal{A}}^{2}C^{2}_{\mathscr{A}}}{2}\|\mathcal{A}^{\frac{1}{2}}u_{m}\|_{H}^{2}+\frac{1}{2}\|\mathcal{A}^{\frac{1+\sigma}{2}}u_{m}\|^{2}_{H},$ (21)
so that after multiplying (2.1.2) by $2$ we are left with
@@
Using (14) with $\frac{1-\sigma}{2}\leq\frac{1+\sigma}{2}$ and Young’s inequality twice we get
@@
Using the estimate (11) on $\mathfrak{B}$, Young’s inequality, the embedding (16) $X\hookrightarrow C([0,T];W)$ and (14) with $\frac{1-\sigma}{2}\leq\frac{\sigma+1}{2}$ we get
$\begin{split}2&\int_{0}^{t}|\langle\mathfrak{B}[u_{m},{u_{*}}]+\mathfrak{B}[{u_{*}},u_{m}],\phi\rangle|ds\\ &\leq 2\int_{0}^{t}C_{\mathfrak{B}}(\|\mathcal{A}^{\frac{1+\sigma}{2}}u_{m}\|_{H}\|\mathcal{A}^{\frac{\sigma}{2}}{u_{*}}\|_{H}+\|\mathcal{A}^{\frac{1+\sigma}{2}}{u_{*}}\|_{H}\|\mathcal{A}^{\frac{\sigma}{2}}u_{m}\|_{H})\|\mathcal{A}^{\frac{1-\sigma}{2}}\phi\|_{H}ds\\ &\leq\int_{0}^{t}\frac{C_{\mathfrak{B}}^{2}}{\|{u_{*}}\|_{X}}(\|\mathcal{A}^{\frac{1+\sigma}{2}}u_{m}\|_{H}^{2}\|\mathcal{A}^{\frac{\sigma}{2}}{u_{*}}\|_{H}^{2}+\|\mathcal{A}^{\frac{1+\sigma}{2}}{u_{*}}\|_{H}^{2}\|\mathcal{A}^{\frac{\sigma}{2}}u_{m}\|_{H}^{2})+{\|{u_{*}}\|_{X}}\|\mathcal{A}^{\frac{1-\sigma}{2}}\phi\|_{H}^{2}ds\\ &\leq{C_{\mathfrak{B}}^{2}}\|{u_{*}}\|_{X}(C_{T}^{2}\int_{0}^{t}\|\mathcal{A}^{\frac{1+\sigma}{2}}u_{m}\|_{H}^{2}ds+\max_{s\in[0,t]}\|\mathcal{A}^{\frac{\sigma}{2}}u_{m}\|_{H}^{2})\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad+(c_{\mathcal{A}}C_{A}+1)^{2}{\|{u_{*}}\|_{X}}\int_{0}^{t}\|\mathcal{A}^{\frac{1+\sigma}{2}}u_{m}\|_{H}^{2}ds\\ &\leq\|{u_{*}}\|_{X}(\max\{C_{T}^{2},1\}C_{\mathfrak{B}}^{2}+(c_{\mathcal{A}}C_{A}+1)^{2})\mathcal{E}[\mathcal{A}^{\frac{\sigma}{2}}u_{m}](t)\end{split}$
@@
${C_{A}}\mathcal{E}[u_{m}](t)+\frac{1}{4}\mathcal{E}[\mathcal{A}^{\frac{\sigma}{2}}u_{m}](t)\leq\delta^{-1}\|f\|_{Y}^{2}+(1+C_{A}c_{\mathcal{A}}^{2})\|u_{m}(0)\|^{2}_{W}.$ (22)
In particular, due to (22), $\|u_{m}\|^{2}_{L^{2}(0,T;U)}\leq C_{U}(\|f\|_{Y}^{2}+\|u_{m}(0)\|^{2}_{W})$ with $C_{U}=4\max\{\delta^{-1},(1+C_{A}c_{\mathcal{A}}^{2})\}$. To estimate $\|\partial_{t}u_{m}\|_{V^{*}}$, we proceed similarly to [Evans:2010, Chapter 7.1]. Let $\phi\in V\subset H$ with $\|\phi\|_{V}\leq 1$. We write $\phi=\phi^{1}+\phi^{2}$ with $\phi^{1}\in U_{m}$ and $\phi^{2}\in U_{m}^{\perp}$, where the orthogonal complement is taken with respect to the inner product of $H$. Since $\partial_{t}u_{m}\in U_{m}\subset H$, we infer from equation (19)
$\begin{split}\langle\partial_{t}u_{m},\phi\rangle_{V}=\langle\partial_{t}u_{m},\phi^{1}\rangle&=\langle f-\mathfrak{A}u_{m}-\mathfrak{B}[u_{m},{u_{*}}]-\mathfrak{B}[{u_{*}},u_{m}],\phi^{1}\rangle\\ &\leq\|f-\mathfrak{A}u_{m}-\mathfrak{B}[u_{m},{u_{*}}]-\mathfrak{B}[{u_{*}},u_{m}]\|_{V^{*}}\|\phi^{1}\|_{V}.\end{split}$
Moreover, we have
$\|\phi^{1}\|_{V}^{2}=\|\mathcal{A}^{\frac{1-\sigma}{2}}\phi^{1}\|_{H}^{2}=\|\sum_{k=1}^{m}\lambda_{k}^{\frac{1-\sigma}{2}}\langle\phi,\phi_{k}\rangle\phi_{k}\|^{2}_{H}=\sum^{m}_{k=1}\lambda_{k}^{1-\sigma}\langle\phi,\phi_{k}\rangle^{2}\leq\|\phi\|_{V}^{2},$
where the last equality follows from Parseval’s theorem. Thus, using (11), (16)
$\begin{split}\|\partial_{t}u_{m}\|_{V^{*}}&\leq\|\mathfrak{A}u_{m}\|_{V^{*}}+\|\mathfrak{B}[u,{u_{*}}]\|_{V^{*}}+\|\mathfrak{B}[{u_{*}},u]\|_{V^{*}}+\|f\|_{V^{*}}\\ &\leq C_{\mathfrak{A}}\|u_{m}\|_{U}+C_{\mathfrak{B}}\|{u_{*}}\|_{X}(\|u_{m}\|_{U}+\|u_{m}\|_{W})+\|f\|_{V^{*}}\\ &\leq C_{V}\|u_{m}\|_{U}+\|f\|_{V^{*}}\end{split}$
@@
$\|\partial_{t}u_{m}\|_{L^{2}(0,T;V^{*})}^{2}\leq 2C_{U}C_{V}^{2}(\|f\|_{Y}^{2}+\|u_{m}(0)\|^{2}_{W})+\|f\|_{Y}^{2}=2(C_{U}C_{V}^{2}+1)(\|f\|_{Y}^{2}+\|u_{m}(0)\|^{2}_{W}).$ (23)
Thus, combing this with (22) we obtain the a priori estimate
$\|u_{m}\|_{X}\leq C_{G}(\|f\|_{Y}+\|u_{m}(0)\|_{W})$ (24)
@@ 2.1.3 Existence and uniqueness
It is therefore left to show that $u(0)=g$ holds, which we do by closely following the argument in [Evans:2010, Section 7.1.2 c]. First of all, since $X\hookrightarrow C([0,T];W)$, we know that imposing an initial condition as $u(0)$ is sensible. Then by partial integration with respect to time in (2.1.3), we have on the one hand for some $v\in C^{1}([0,T];W)\subset L^{2}(0,T;V)$ with $v(T)=0$
$\langle\partial_{t}v,u\rangle_{Y}+\langle\mathfrak{A}u+\mathfrak{B}[u,{u_{*}}]+\mathfrak{B}[{u_{*}},u],v\rangle_{Y}=\langle f,v\rangle_{Y}+\langle u(0),v(0)\rangle_{V^{*}}$ (25)
@@
$\lim_{n\to\infty}-\langle\partial_{t}v,u_{m}\rangle_{Y}+\langle\mathfrak{A}u_{m}+\mathfrak{B}[u_{m},{u_{*}}]+\mathfrak{B}[{u_{*}},u_{m}],v\rangle_{Y}=\langle f,v\rangle_{Y}+\langle g,v(0)\rangle_{V^{*}}$ (26)
since $u_{m}(0)\to g$ weakly in $V$. Comparing both expression we see that due to arbitrariness of $v(0)$, we obtain $u(0)=g$. Uniqueness of $u$ follows by standard arguments, due to the linearity of (2), see [QuadraticWave, Section 3.2.3] for details. ∎
@@ 2.2 Contraction argument for the bilinear problem
$\|\mathcal{F}^{\prime}_{u_{1}}-\mathcal{F}^{\prime}_{u_{2}}\|_{\mathcal{L}(X,Y)}\leq K\|u_{1}-u_{2}\|_{X},\quad u_{1},\,u_{2}\in D_{*}.$ (27)
@@
A proof of Lemma 2.1 can be found in e.g. [Kantorovich1948, KantorovichAkilov]. With this, for any given $T>0$ we can establish well-posedness of the abstract system, provided the data are sufficiently small.
@@
Let $T>0$, $(f,g)\in L^{2}(0,T;D(\mathcal{A}^{\frac{1-\sigma}{2}})^{*})\times{D}(\mathcal{A}^{\frac{\sigma}{2}})$ and Assumption 2.1 hold with $\sigma\in[0,1]$. Then, there exists a $r>0$ so that if
@@
The proof follows a similar argumentation as the proof of [QuadraticWave, Theorem 3.2]. For convenience of the reader, we restate it here. As in Theorem 2.1, we use the short notations $Y:=L^{2}(0,T;D(\mathcal{A}^{\frac{1-\sigma}{2}})^{*})$, $U:=D(\mathcal{A}^{\frac{1+\sigma}{2}})$, $V:=D(\mathcal{A}^{\frac{1-\sigma}{2}})$ and $W:={D}(\mathcal{A}^{\frac{\sigma}{2}})$
@@
$\mathcal{F}u:=(\partial_{t}u+\mathfrak{A}u+\mathfrak{B}[u,u],u(0))-(f,g).$ (31)
@@
Then, by assumptions 2.1(iii) (11) and (16) that yield boundedness of $\mathfrak{B}:X\times X\to Y$ we have
$\frac{\|\mathfrak{B}[h,h]\|_{Y}}{\|h\|_{X}}\leq\frac{(C_{\mathfrak{B}}^{2}C_{T}^{2}\|h\|_{X}^{2}\int_{0}^{T}\|h\|_{U}^{2}ds)^{\frac{1}{2}}}{\|h\|_{X}}\leq{C_{\mathfrak{B}}}C_{T}\|h\|_{X}\to 0\quad(\|h\|_{X}\to 0)$ (32)
@@
$\mathcal{F}^{\prime}_{u_{*}}:\xi\mapsto(\partial_{t}\xi+\mathfrak{A}\xi,\xi(0))+(\mathfrak{B}[\xi,{u_{*}}]+\mathfrak{B}[u_{*},\xi],0),$ (33)
@@
$\|(\mathcal{F}^{\prime}_{u_{1}}-\mathcal{F}^{\prime}_{u_{2}})u\|_{Y}\leq\|\mathfrak{B}[u,u_{1}-u_{2}]+\mathfrak{B}[u_{1}-u_{2},u]\|_{Y}\leq 2C_{\mathfrak{B}}C_{T}\|u_{1}-u_{2}\|_{X}\|u\|_{X}$ (34)
@@
$\beta K\|-(\partial_{t}\cdot+\mathfrak{A}\cdot,\cdot(0))^{-1}(f,g)\|_{X}\leq\frac{1}{2}.$ (35)
@@
We note that the above theorem does not provide a global well-posedness result, as the choice of $r$ depends on the embedding constant $C_{T}$, which can depend on $T$.